OpenStudy (anonymous):

You invest $100,000 in an account with an annual interest rate of 4.5%, compounded semiannually. How much money is in the account after 15 years? Round your answer to the nearest whole number.

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi :P

OpenStudy (misty1212):

\[100,000\times \left(1+\frac{.045}{2}\right)^{30}\] and a calculator should do it

OpenStudy (misty1212):

semi annually means twice a year, i used the formula with \(n=2\)

OpenStudy (anonymous):

okay thank you so much!!